OpenStudy (anonymous):

r(t)= (1-t)<-2,-1> + t<1,2> I don't know how to get to the answer which I know is this. =<-2+3t, -1+3t> This is parameterization for integral of a line segment and this part is holding me up. This is the example I could find but I'm not sure where they got the vector.

OpenStudy (anonymous):

\[r \left( t \right)=(1-t)<-2,-1>+t<1,2>=(1-t)(-2 i-1 j)+t(1 i+2 j)\] \[=-2 i-1j+2ti+t j+t i+2 t j=(-2+2 t+t) i+(-1+t+2 t) j=(-2+3 t)i+(-1+3t) j\] =<-2+3t,-1+3t>

OpenStudy (anonymous):

Well, I obviously need a break today. Thanks! Very helpful.

OpenStudy (anonymous):

yw